Pls. help me regarding this example. I'll be taking a test of these soon. I just want to know how this is solved, I'm having trouble with the reversed word. Thanks.
The sum of the digit of a two no. is 9. If the digits are reversed, the numbers is increased by 27. Find the original no.

Answer by Alan3354(69443)

The sum of the digit of a two no. is 9. If the digits are reversed, the numbers is increased by 27. Find the original no.
----------
It's confusing, because we don't think of a 2 digit # as 10 x the left digit plus the right digit.
-----------------
A 2 digit number is xy.
x+y = 9
The value of xy is 10x + y.
Reversed, 10y + x = 10x + y + 27
-9x + 9y = 27
-x + y = 3
x + y = 9
---------- Add
2y = 12
y = 6
x = 3
---------
Check:
36 = 63 - 27, so it's correct.